OpenStudy (anonymous):

Which ordered pair does not satisfy the system? y <-2x + 9 and y > 3/4 x - 2 (3, -2) (2, 3) (-1, 4) (-1, -1

OpenStudy (rulnick):

(3,-2)

OpenStudy (rulnick):

hold on

OpenStudy (rulnick):

yes, confirmed see second relation: y > (3/4)x-2 at (x,y)=(3,-2) this is -2 > (3/4)(3)-2 = 9/4-2=1/4 false so we're good, it's (3,-2) thanks for your patience while I verified

OpenStudy (anonymous):

Ok Thank You!